OpenStudy (anonymous):

Can you help me understand how to do this problem. Given the function f(x) = 4(x+3) - 5, solve for the inverse function when x = 3. (1 point)

OpenStudy (anonymous):

To make this easier, let y = f(x). So, we have: \[y = 4(x + 3) - 5\]To solve for the inverse, switch x and y. \[x = 4(y + 3) - 5\]Now, solve for y again in terms for x. Can you take it from there?

OpenStudy (anonymous):

3=4y+12 - 5 so 3=4y+7 so -4=4y so y=-1?

OpenStudy (anonymous):

Yes! Good job! Also, there's a cool little trick to inverse functions like this. If you're trying to find what the inverse is at an x-value, you can just plug that in for y (or f(x)) in the original and solve for x.

OpenStudy (anonymous):

Can you give me an example? Im not sure I understand that.

OpenStudy (anonymous):

what about this when you don't have a value for x and need to do the inverse? Given the function f(x) = 5x-3/4 , which of the below expressions is correct? f-1(x) = 3-5x/4 f-1(x) = 4x+3/5 f-1(x) = 4x-3/5 f-1(x) = -5x-3/4

OpenStudy (anonymous):

Follow the same steps as in the first example, but this time, since you don't have any values to plug in, leave the x there and solve for y in terms of x. Then, once you get y by itself, replace it with \(f^{-1}(x)\).

OpenStudy (anonymous):

So, to get you started, \[f(x) = 5x-\frac34\]\[y=5x-\frac34\]\[x=5y-\frac34\]

OpenStudy (anonymous):

it's actually 5x-3 over 4

OpenStudy (anonymous):

it wouldn't copy the equation correctly.

OpenStudy (anonymous):

Ah ok. No problem. \[f(x) = \frac{5x-3}{4}\]\[y=\frac{5x-3}{4}\]\[x=\frac{5y-3}{4}\]

OpenStudy (anonymous):

then how do i do the f-1(x)

OpenStudy (usukidoll):

split the fractions up \[x = \frac{5y}{4}- \frac{3}{4}\]

OpenStudy (usukidoll):

now add 3/4 to both sides.

OpenStudy (anonymous):

Well, remember when we let y = f(x)? Solve for y first, then you can just replace the y with the \(f^{-1}(x)\). It's just a notation thing. You can work with f(x) the entire time, it's just a bit messy.

OpenStudy (anonymous):

Do you need help with the algebra to solve for y or can you do that part?

OpenStudy (anonymous):

can you help? I got y=3+4X? Is that right?

OpenStudy (usukidoll):

let me check that for a sec

OpenStudy (anonymous):

Not quite. I think you forgot the 5 somewhere.

OpenStudy (usukidoll):

\[x + \frac{3}{4}= \frac{5y}{4}\] then divide 5/4 throughout the entire equation .

OpenStudy (anonymous):

\[x = \frac{5y - 3}{4}\]\[4x = 5y - 3\]\[5y = 4x + 3\]It looks like you got to here in your algebra. All that's left to do is divide by 5 to get y by itself. Do you get it?

OpenStudy (usukidoll):